Please give 10 examples of "The Product of the Form".
Examples:
1.(x+y)(x²-xy+y²) = x³+y³
2.(2x+2y)(2x²-2xy+2y²) = 4x³+4y³
Just like that. So, please help me. Please answer my question immediately. Thanks in advance! God bless. ;)

Answer by lawdej(58)


you can produce so many examples based on the expression (x+y)(x²-xy+y²) = x³+y³ or(x-y)(x²+xy+y²) = x³-y³ but first let me correct the second example you put b/c it is wrong.here is the correct one
2.(2x+2y)((2x)²-4xy+(2y)²) = (2x)³+(2y)³= 8x³+8y³
additional examples
1.(x+2)(x²-2x+2²) = x³+2³
2.(6-3)(6²+18+3²) = 6³-3³
3.(2x+5)((2x)²-10x+5²) = (2x)³+5³
4.(3+2y)(3²-6y+(2y)²) = 3³+(2y)³
5.(4x+4y)((4x)²-16xy+(4y)²) = (4x)³+(4y)³
6.(3+2)(3²-6+2²) = 3³+2³
7.(x+2y)(x²-2xy+(2y)²) = x³+(2y)³
8.(5x+6y)((5x)²-30xy+(6y)²) = (5x)³+(6y)³
9.(10x+5y)((10x)²-50xy+(5y)²) = (10x)³+(5y)³
10.(4x-4y)((4x)²+16xy+(4y)²) = (4x)³-(4y)³
11.(-2x+4y)((4y)²+8xy+(2x)²) = (4y)³-(2x)³ etc....